Abby13915:

One gallon of paint usually covers about 350 square feet. How many square feet will 2.5 gallons cover? Show your work, including labeling your answer with the appropriate units of measure.

NemoEmo:

350*2 = 700 350/2=175 700+175=875

DontQuestionMe:

@nemoemo wrote:
350*2 = 700 350/2=175 700+175=875
Dont forget the units so it would be 875 square feet covered

NemoEmo:

2.5 gallons can cover 875 square feet

NemoEmo:

@dontquestionme wrote:
@nemoemo wrote:
350*2 = 700 350/2=175 700+175=875
Dont forget the units so it would be 875 square feet covered
yeah I forgot

DontQuestionMe:

@nemoemo wrote:
@dontquestionme wrote:
@nemoemo wrote:
350*2 = 700 350/2=175 700+175=875
Dont forget the units so it would be 875 square feet covered
yeah I forgot
Its okk

Abby13915:

thank you